OpenStudy (chris215):

If f(1) = 5 and f '(x) 6/x^2+2 , which of the following is the best approximation for f(1.03) using local linearization?

jimthompson5910 (jim_thompson5910):

is this what f ' (x) is? \[\Large f \ ' (x) = \frac{6}{x^2+2}\] right? or no?

OpenStudy (chris215):

yeah

jimthompson5910 (jim_thompson5910):

ok what is f ' (1) equal to?

OpenStudy (chris215):

2?

jimthompson5910 (jim_thompson5910):

yes, so m = 2 is the slope of the tangent line

jimthompson5910 (jim_thompson5910):

f(1) = 5 means we know (1,5) is on f(x) so x = 1 and y = 5 y = mx+b y =2x + b ... plug in m = 2 5 = 2*1 + b ... plug in (x,y) = (1,5) ... ... ... b = ???

OpenStudy (chris215):

so 3?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

L(x) = 2x + 3 is the linear approximating line at (1,5) so to approximate f(1.03) you need to calculate L(1.03)

OpenStudy (chris215):

i got 5.05

jimthompson5910 (jim_thompson5910):

2*1.03+3 = 5.06

OpenStudy (chris215):

thak you so much !

jimthompson5910 (jim_thompson5910):

you're welcome